GMT Date: 073
Calendar Date: 03/14/03
Payloads currently on-board ISS:
BSTC, CGBA, CSLM-2, InSPACE, MAMS, MSG (facility), PFMI, SAMS, STES-10, SUBSA and ZCG.
MAMS, SAMS and STES-10 continue nominal, basically unattended, operations. BSTC, CGBA, CSLM-2, InSpace, PFMI, SUBSA and ZCG are inactive. MSG is in a trouble-shooting mode.
EXPRESS RACKS 1 and 4 (ER1 & ER4) are powered and supporting payload operations. ER2, ER3 & ER5 are currently powered down.
The MSG circuit breaker problem evaluation continues, and troubleshooting procedures are being implemented carefully. Thirteen sensor cables were disconnected from MSG on day 59, leaving five sensors connected that were verified to cause no problem with MSG powered. The disconnected cables were reconnected one at a time to attempt isolate the connection that is causing an MSG circuit breaker to trip. Prior to today, twelve cables were reconnected and powered with no apparent anomalous results. Reconnection of the thirteenth cable was performed today with no apparent anomalous result. The planned fan replacement activity was also performed, and the new fan appears to be operating nominally. MSG was power cycled by the MSG Ops Team, and data was collected for further analysis. OCR msg1all000077 has been submitted to prepare for continuation of troubleshooting steps on Monday, GMT day 076. This OCR calls for power cycling ESEM 4A outlet 2 (MLC), ESEM 4B outlet 2, and ESEM 4A outlet 1 (Video drawer) sequentially, allowing each outlet an hour of operation.
InSpace Coil assemblies may need to be inspected again prior to beginning test operations if it becomes feasible to install the payload in the MSG.
MAMS is operating nominally in the quasi-steady mode. HiRAP is currently inactive. MAMS data is available on the acceleration measurement web page.
PFMI hardware removal from the MSG is contingent upon success of MSG trouble-shooting.
All SAMS sensors in EXPRESS RACKS are operating nominally. The SAMS sensor in the MSG is currently turned-off. SAMS data is also available on the acceleration measurement web page.
STES-10 health and status data is nominal. STES-10 is configured with diffusion-controlled chambers that require no crew operation activities.
Week 15 (3/10-3/16) - Planning for this week was based on approximately 10 hours of crew time for NASA payload operations. The crew will continue preparations for a newly planned EVA during week 17. A ground commanded reboost of the ISS was performed yesterday, and another one is planned for later today. MSG troubleshooting occurred every day this week, and will continue next week.
Week 16 (3/17-3/23) – Preliminary planning for this week is being based on approximately 22 hours of crew time for NASA payload operations. The flight crew will continue preparations for the newly planned EVA during week 17. Time is being allocated on four days of this week to continue MSG trouble-shooting.
M. Vlasse, PhD, DSc
PSR Increment Scientist
BCSS Biotechnology Cell Science Stowage
BSTC Biotechnology Specimen Temperature Controller (General status information on this payload will include coverage of the following items of support equipment: BTR, BCSS and GSM)
BTR Biotechnology Refrigerator
CGBA Commercial Generic Bioprocessing Apparatus
ER1 EXPRESS RACK 1
ER2 EXPRESS RACK 2
ER3 EXPRESS RACK 3
ER4 EXPRESS RACK 4
ER5 EXPRESS RACK 5
GSM Gas Supply Module
InSPACE Investigating the Structure of Paramagnetic Aggregates
MAMS Microgravity Acceleration Measurement System
MSG Microgravity Science Glovebox
OCR Operations Change Request
PAO Public Affairs Office
PFMI Pore Formation and Mobility Investigation
PGB Plant Generic Bioprocessing Apparatus
PSR NASA Code U Physical Sciences Research Division
RIC Rack Interface Controller
SAMS Space Acceleration Measurement System II
SPD NASA Code U Space Products Development Division
STES Protein Crystal Growth - Single Thermal Enclosure System
SUBSA Solidification Using a Baffle in Sealed Ampoules
ZCG Zeolite Crystal Growth